# Let It Flow: Simultaneous Optimization of 3D Flow and Object Clustering

### v1

@@ 3.3 Hard rigidity loss
3.4 Soft rigidity loss
Intuitively, when the cluster $C$ is incorrect (e.g. it overflows into neighbouring objects), the optimization of the flow through the hard rigid loss delivers inaccurate flow. In order to enable outlier rejection, we introduce $k$-dimensional soft clustering vector $\mathbf{v}\in\mathopen{[}0,1\mathclose{]}^{(k+1)},\|\mathbf{v}\|=1$ that is supposed to softly selects high-reward edges. This vector models how much the points are likely to be in the rigid object that is dominant within the cluster. Given this notation, we define the soft rigidity score induced by the point $\mathbf{p}_{m}$ as
@@ 3.4 Soft rigidity loss
4 Experiments
Datasets.
We conducted experiments using standard scene flow benchmarks. Firstly, we employed two large-scale, well-known LiDAR autonomous driving datasets: Argoverse1 [3] and also newer version Argoverse2 [40] and Waymo [34] open dataset. These datasets encompass challenging dynamic scenes with various maneuvers captured by different LiDAR sensor suites. To ensure a fair comparison, we sampled and processed the LiDAR datasets according to the methodology outlined in [20] and use validation split sequences first frames. As there are no official scene flow annotations, we adopted the data processing approach from [15] to derive pseudo-ground-truth scene flow information based on object detection annotations, akin to the methodology employed in [20]. Finally, we removed ground points in the height of 0.3 meters or lower for Waymo dataset, following the procedure outlined in previous works [24, 20, 17, 21, 11, 36]. For Argoverse1 and Argoverse2, we used the accompanying ground maps to remove ground points and constrain the range of input points to 35 meters as done in [7, 11, 41, 17].
Subsequently, we utilized the stereoKITTI dataset [27, 24], which comprises real-world scenes from autonomous driving scenarios. The point clouds in the stereoKITTI dataset are however constructed by lifting the depth from stereo images by calculated optical flow, resulting in 3D points with one-to-one correspondence and huge amount of dynamic points per object [7]. The dataset was further divided by [28] into a testing part referred to as KITTIt.
Implementation Details.
We use $k=16$ and set $\alpha$ and $\beta$ loss weights to 1. We use the learning rate of 0.004. The $\theta$ distance threshold for outlier rejection is set to 0.03 as in [5, 36]. We perform hard clustering by spatio-temporal DBSCAN over horizon of 5 point clouds and set epsilon parameter to 0.3 and minimal samples to 1, resulting in clustering the points to the same cluster if the Euclidian distance is lower than 0.3. The flows are optimized until a fixed number of iterations (1500) or until reaching the convergence of the loss function. For other comparative methods, we use the same optimized parameters as reported in their original papers and official implementations.
Evaluation metrics.
For proper evaluation of results we need to define point error $e_{i}$ and relative point error $e_{i}^{r}$. Following the notation introduced in Section 3, we define these errors in meters as follows:
@@ 4 Experiments
In automotive scenes, the majority of the points come from a static background. When the metrics ($EPE$, $AS$, $AR$) are calculated and averaged for the full point cloud, the results show mostly performance on static points, lowering the importance of dynamic objects [7, 35]. If we dig even further, the dynamic points are heavily dominated by large vehicles and the performance on smaller, yet equally important classes, like Pedestrians and Cyclists are not observable from the dynamic $EPE$ metrics. Therefore, we also report the metrics per-class on the main benchmark, Argoverse2.
4.1 Comparison with State-of-the-Art Methods
Table 1: Performance on Argoverse2 dataset in foreground aware metrics, i.e. Threeway EPE metric.
@@ 4.1 Comparison with State-of-the-Art Methods
Table 2: Performance on Argoverse2 dataset in object class aware EPE.
@@
Methods Avg.normal-$\downarrow$ Dyn.normal-$\downarrow$ Stat.normal-$\downarrow$ Avg.normal-$\downarrow$ Dyn.normal-$\downarrow$ Stat.normal-$\downarrow$ Avg.normal-$\downarrow$ Dyn.normal-$\downarrow$ Stat.normal-$\downarrow$
@@
We benchmark our method against the top-performing methods, such as NSFP [20], MBNSFP [36] and Chodosh [14], which all share the same structural optimization-based regularization, i.e Neural Prior [20, 21]. For all methods, we use the official implementation provided in papers. For Chodosh [7], we used the codebase published in [35]. For each method, we compensate the ego-motion with KISS-ICP [37] first, then estimate the flow by methods.
@@
In Table 1, we show results on Argoverse2, the main benchmark for 3D scene flow estimation. Our method dominates the dynamic foreground metrics, while the MBNSFP [36] has better static foreground. We suggest, that their reliance on Neural prior architecture coupled with under-segmented rigidity tends to overfit on the static classes, since they share a single movement, caused by ego-motion, and the MLP layers in Neural Prior do not have expressive capacity to catch multiple motion patterns. Chodosh [7] do not optimize rigidity and Neural prior jointly, but fine-tune the rigidity as a post-processing step, resulting in more separable object motion. Our method without Neural Prior and merging clusters achieves the best average EPE over the dynamic and static metrics.
@@
When zooming in to per-class scene flows, we observe a trend of overfitting to larger dynamic objects, i.e. vehicles. We suggest, that fitting one structural neural prior to multiple objects in the scene leads to the local minima of solving the objects with the most of points, and sacrificing the smaller ones. In Table 2, we show the performance on Pedestrians, Cyclists and Vehicles separately. For example, we can see, that the performance of the other methods on Pedestrian class is halved, compared to ours, where the over-segmentation of the scene usually safely cluster the whole Pedestrian as one hard cluster and estimate rigid flow without structural neural constraints. We see that in Cyclist category, the performance of ours is even higher. We do not lose the ability estimate flows on Vehicles, as we can separate them in our merging clustering as well. We have worse static vehicles compared to MBNSFP[36], which we explain by overfitting into static flow.
4.2 Neural Prior with Hard and Soft Clusters.
@@
In order to compare our proposed rigidity via hard and soft clusters to the most similar methods (MBNSFP [36] as it also use spatial consistency with outlier rejection), we change the direct optimization of flows for neural prior architecture as in other methods [20, 36, 7]. We also perform experiments under the metrics proposed in MBNSFP [36], i.e. overall end point error without foreground/background split.
@@
In Table 3, we see the results on Waymo [34] and Argoverse1 [3] datasets on splits used in  [36]. We can see, that the better performance is reached with our rigidity regularization. The MBNSFP [36] is the second top-performing methods with regular Neural prior [20] with cycle consistency and structural regularization behind.
Runtime Benefit of Implementing Soft Clusters as Neighborhoods.
We analyze the performance-time trade-off in Figure 3 by acquiring the $EPE$ based on inference time for our method on Argoverse1 dataset while using Neural Prior architecture with proposed Soft Clusters. The speed was measured on NVIDIA Tesla V100 GPU on complete point clouds. We compare with optimization-based methods FastNSF [21] and MBNSF [36] with distance transform acceleration proposed by [21] and DBSCAN clustering, and Neural prior [20] without distance transform acceleration [21].
@@
Table 3: Performance on LiDAR datasets. We show performance on standard LiDAR benchmarks with overall metrics (without dynamic/class split). Results are averages over 3 runs.
@@
Compared to the MBNSFP [36], our method is approximately $\mathbf{10}\times$ faster while stopping at the same prediction error. Even though their spatial consistency regularization uses the same algorithm of computing the point-to-point displacements with outlier rejection, our method is significantly faster because we design the clusters on the same-sized neighborhoods, which can be efficiently parallelized on GPUs. In contrast to our method, MBNSFP computes eigen vectors of matrix with variable-sized clusters [9]. Thus, MBNSFP is looping over $c$ clusters $\{N_{0},N_{1},...,N_{c}\}$ one-by-one, whereas our method allows GPU to parallel the $\mathbb{R}^{N\times k\times k}$ tensor computation for $N$ points in point cloud and $k$ neighbors simultaneously.
Figure 3: Flow estimation error vs. run time on Argoverse1 dataset. Our method in green offers speed-accuracy trade-off that outperforms the converged baselines on overall EPE metric.
4.3 Results on StereoKITTI benchmark.
@@
Our main baselines for performance comparison are state of the art self-supervised methods such as SCOOP [17], SLIM [1], FastNSF [21], Rigid Flow [19]. Our method achieves the new self-supervised state-of-the-art performance on stereoKITTI benchmark with the relative improvement of 38% of $EPE$ compared to top-performing SCOOP [17] on StereoKITTI and also achieves the improvement of 42% of $EPE$ compared to the top-performing NSFP [21] on KITTIt.
Table 4: Comparison with State-of-the-Art on KITTI dataset. We evaluate scene flow based on standard metrics $EPE$, $AS$, $AR$ and $Out.$ for StereoKITTI dataset. Previous state of the art performance is underlined and current best is in bold. Our results averaged over 5 runs.
Test data Method $EPE(m)$downarrow $AS(\%)$uparrow $AR(\%)$uparrow $Out.(\%)$downarrow
@@
4.4 Ablation Study
Method components
In Table 5, we show metrics when adding one enhancement per time. We observe the benefits of components on final performance on the Argoverse2 dataset. Without rigidity regularization and clusters a), the methods resembles flow only to the nearest neighbors in target point cloud and therefore fails to produce meaningful motions. By including the hard cluster rigidity b), we observe physically consistent motions. However, since we enforce rigidity on over-segmented clusters, the objects still deforms. With addition of soft clusters c), we are able to connect nearby points from different hard segments to rigidity term and reject outliers, expanding the object rigidity. By guiding the hard clusterization with flows d), we acquire small boost by connecting lonely clusters divided by occlusions or separate by LiDAR sampling rate back to the major rigid body.
Table 5: Ablation study for individual method components on Argoverse2.
Module Avg. $EPE(m)$downarrow Dyn. Fore. $EPE(m)$downarrow Stat. Fore. $EPE(m)$downarrow Stat. Back. $EPE(m)$downarrow
@@ 4.4 Ablation Study
5 Limitations
@@
Even though our method is reasonably fast, it still does not, as the other methods, achieve real-time performance while keeping high accuracy. Potential direction for improvement is shown in [35]. Secondly, our method does not deal with the local minima of the distance loss, resulting in sub-optimal flows even when the object is correctly clustered. These two issues are worth of the future research to improve scene flow further.
6 Conclusion
@@
We presented a novel self-supervised method for 3D scene flow prediction inspired by rigid motion of multiple objects in the scene. The method introduced overlapping cluster rigidity regularization, that achieve state-of-the-art on standard benchmarks and is tuned towards the complex dynamic objects rather then static backgrounds. The framework is part of the optimization-based models family a therefore does not require any training data. The proposed solution cleverly adapts rigidity mechanism to discover long tail object flows showing that it is important, how are the points paired into the regularization mechanism.
Appendix 0.A Datasets and Code Implementation
For datasets splits in comparison with MBNSFP [36], we maintain the philosophy of the aforementioned method and extract first frames from each sequence of Argoverse [3] and Waymo [34]. For the main experiments against [7, 20, 36], we extract the third frame from each Argoverse2 [3] sequence, as we perform spatio-temporal clustering and past frames are always available. That resembles 150 unique scenes with various agents, motion and ego manuevers including rotational movement. We apply KISS-ICP [37] on the sequences to compensate ego-motion for all methods.
We used implementation of Chodosh [7] from https://github.com/kylevedder/zeroflow, as it comes from the same group and no other official implementation is available. For MBNSFP [36], we used the official implementation from https://github.com/kavisha725/MBNSF/tree/main.
Our implementation is available on https://github.com/vacany/let-it-flow
Appendix 0.B DBSCAN Clustering in Scene Flow
In order to achieve the best possible clustering in competing methods, the authors of [36, 7] tune The DBSCAN algorithm to catch the bigger objects, which is visible in parameter epsilon (distance grouping threshold) set to 0.8 and minimal number of point samples in one density-based cluster set to 30 in MBNSFP [36]. Such parametrization allows for successful grouping of points in whole bodies of large objects such as vehicles, but simultaneously group multiple smaller objects together. Another issue stems of points, that falls below such threshold and therefore do not form the cluster and are treated without rigidity regularization in the methods. See example of such DBSCAN clustering in Figure 5, where the multiple pedestrians are grouped together with a tree or a vehicle in the middle of point cloud has blue points (points not assigned to any cluster).
We conclude that such clustering parametrization leads to over-fiting on large objects a therefore we designed our hard clustering approach to assign all points to clusters in a "over-segmentation" manner and gradually merge the over-segmented clusters based on flow optimization.
Figure 5: Qualitative example on Waymo Dataset. We show, that by DBSCAN under-segmentation (left part of the image, where color denotes cluster id), it is impossible to separate motions of multiple objects in one cluster. When the MBNSFP [36] method converges, the flow vectors on a single cluster looks uniform, resembling similar motion patterns on separate objects, while with our gradual clustering, we are able to separate motions and assign correct headings of flows. In the last row, we applied DBSCAN clustering to only flow vectors to show, that we can separate the objects based on our method output flows. We could not find parameters to separate objects with MBNSFP [36] flows, as they are too close to each other.

### v2

@@ III-C Hard rigidity loss
III-D Soft rigidity loss
Intuitively, when the cluster $C$ is incorrect (e.g., it overflows into neighboring objects), the optimization of the flow through the hard rigid loss delivers inaccurate flow. In order to enable outlier rejection, we introduce $k$-dimensional soft clustering vector $\mathbf{v}\in\mathopen{[}0,1\mathclose{]}^{(k+1)},\|\mathbf{v}\|=1$ that is supposed to softly selects high-reward edges. This vector models how much the points are likely to be in the rigid object that is dominant within the cluster. Given this notation, we define the soft rigidity score induced by the point $\mathbf{p}_{m}$ as
@@ III-D Soft rigidity loss
IV Experiments
Datasets
We conducted experiments using standard scene flow benchmarks. Firstly, we employed two large-scale, well-known LiDAR autonomous driving datasets: Argoverse1 [3] and the newer version Argoverse2 [41] and Waymo [35] open dataset. These datasets encompass challenging dynamic scenes with various maneuvers captured by different LiDAR sensor suites. To ensure a fair comparison, we sampled and processed the LiDAR datasets according to the methodology outlined in [20] and used the first frames of all available validation sequences as done in [20] for benchmarking. As there are no official scene flow annotations, we adopted the data processing approach from [15] to derive pseudo-ground-truth scene flow information based on object detection annotations, akin to the methodology employed in [20]. Finally, we removed ground points in the height of 0.3 meters or lower for Waymo dataset, following the procedure outlined in previous works [25, 20, 17, 21, 11, 37]. For Argoverse1 and Argoverse2, we used the accompanying ground maps to remove ground points and constrain the range of input points to 35 meters as done in [7, 11, 42, 17].
Subsequently, we utilized the stereoKITTI dataset [28, 25], which comprises real-world scenes from autonomous driving scenarios. The point clouds in the stereoKITTI dataset are, however, constructed by lifting the depth from stereo images by calculated optical flow, resulting in 3D points with one-to-one correspondence and a huge amount of dynamic points per object [7]. The dataset was further divided by [29] into a testing part referred to as KITTIt.
Implementation Details
We use $k=16$ and set $\alpha$ and $\beta$ loss weights to 1. We use the learning rate of 0.004. The $\theta$ distance threshold for outlier rejection is set to 0.03 as in [5, 37]. We perform hard clustering by spatio-temporal DBSCAN over the horizon of 5-point clouds and set the epsilon parameter to 0.3 and minimal samples to 1, resulting in clustering the points to the same cluster if the Euclidian distance is lower than 0.3. The flows are optimized until a fixed number of iterations (1500) or until the convergence of the loss function is reached. For other comparative methods, we use the same optimized parameters as reported in their original papers and official implementations.
Evaluation metrics
For proper evaluation of results, we need to define point error $e_{i}$ and relative point error $e_{i}^{r}$. Following the notation introduced in Section III, we define these errors in meters as follows:
@@ IV Experiments
In automotive scenes, the majority of the points come from a static background. When the metrics (EPE, AS, AR) are calculated and averaged for the full point cloud, the results show mostly performance on static points, lowering the importance of dynamic objects [7, 36]. If we dig even further, the dynamic points are heavily dominated by large vehicles, and the performance on smaller yet equally important classes, like Pedestrians and Cyclists, are not observable from the dynamic EPE metrics. Therefore, we also report the metrics per class on the main benchmark, Argoverse2 [41].
IV-A Comparison with State-of-the-Art Methods
TABLE I: Performance on Argoverse2 dataset in foreground aware metrics, i.e. Threeway EPE metric. Methods sorted by Threeway EPE Average.
@@ IV-A Comparison with State-of-the-Art Methods
TABLE II: Performance on Argoverse2 dataset on object-class aware EPE.
@@
Methods Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$ Avg.$\downarrow$ Dyn.$\downarrow$ Stat.$\downarrow$
@@
We benchmark our method against the top-performing methods, such as NSFP [20], MBNSFP [37] and Chodosh [14], which all share the same structural optimization-based regularization, i.e Neural Prior [20, 21]. For all methods, we use the official implementation provided in the papers. For Chodosh [7], we used the codebase published in [36]. For each method, we first compensate the ego-motion with KISS-ICP [38], then estimate the flow by methods.
@@
In Table I, we show results on Argoverse2, the main benchmark for 3D scene flow estimation. Our method dominates the dynamic foreground metrics, while the MBNSFP [37] has a better static foreground. We suggest that their reliance on Neural prior architecture coupled with under-segmented rigidity tends to overfit on the static classes since they share a single movement caused by ego-motion, and the MLP layers in Neural Prior do not have the expressive capacity to catch multiple motion patterns. Chodosh [7] does not optimize rigidity and Neural prior jointly but fine-tunes the rigidity as a post-processing step, resulting in more separable object motion. Our method without Neural Prior and merging clusters achieves the best average EPE over the dynamic and static metrics.
@@
When zooming in on per-class scene flows, we observe a trend of overfitting to larger dynamic objects, i.e., vehicles. We suggest that fitting one structural neural prior to multiple objects in the scene leads to the local minima of solving the objects with the most points and sacrificing the smaller ones. In Table II, we show the performance on Pedestrians, Cyclists and Vehicles separately. For example, we can see that the performance of the other methods on the Pedestrian class is halved, compared to ours, where the over-segmentation of the scene usually safely clusters the whole Pedestrian as one hard cluster and estimates rigid flow without structural neural constraints. We see that in the Cyclist category, our performance is even higher. We do not lose the ability to estimate flows on Vehicles, as we can separate them in our merging clustering as well. We have worse static vehicles compared to MBNSFP[37], which we explain by overfitting into static flow.
IV-B Neural Prior with Hard and Soft Clusters
@@
To compare our proposed rigidity via hard and soft clusters to the most similar methods (MBNSFP [37] as it also uses spatial consistency with outlier rejection), we change the direct optimization of flows for neural prior architecture as in other methods [20, 37, 7]. We also perform experiments under the metrics proposed in MBNSFP [37], i.e., overall endpoint error without foreground/background split.
@@
In Table III, we see the results on Waymo [35] and Argoverse1 [3] datasets on splits used in  [37]. We can see that better performance is achieved with our rigidity regularization. The MBNSFP [37] is the second top-performing method with regular Neural prior [20] with cycle consistency and structural regularization behind.
@@
TABLE III: Performance on LiDAR datasets. We show performance on standard LiDAR benchmarks with overall metrics (without dynamic/class split). Results are averages over 3 runs.
@@
IV-C Results on StereoKITTI benchmark
@@
Our main baselines for performance comparison are state-of-the-art self-supervised methods such as SCOOP [17], SLIM [1], FastNSF [21], Rigid Flow [19]. Our method achieves the new self-supervised state-of-the-art performance on stereoKITTI benchmark with the relative improvement of 38% of $EPE$ compared to top-performing SCOOP [17] on StereoKITTI and also achieves the improvement of 42% of $EPE$ compared to the top-performing NSFP [21] on KITTIt.
TABLE IV: Comparison with self-supervised State-of-the-Art methods on StereoKITTI dataset. We evaluate scene flow based on standard metrics EPE, AS, AR, and Out. used in StereoKITTI benchmarking. The previous state-of-the-art performance is underlined, and the current best is in bold. Our results averaged over five runs.
[…]
@@
Figure 4: Qualitative example on Waymo Dataset. We show that by DBSCAN under-segmentation (left part of the image, where color denotes assignment to the cluster), it is impossible to separate motions of multiple objects in one cluster. When the MBNSFP [37] method converges, the flow vectors on a single cluster look uniform, resembling similar motion patterns on separate objects, while with our gradual clustering, we are able to separate motions and assign correct headings of flows. In the last row, we applied DBSCAN clustering to only flow vectors to show that we can separate the objects based on our method output flows. We could not find parameters to separate objects with MBNSFP [37] flows, as they are too close to each other.
IV-D DBSCAN Clustering in Scene Flow
In order to achieve the best possible clustering in competing methods, the authors of [37, 7] tune The DBSCAN algorithm to catch the bigger objects, which is visible in parameter epsilon (distance grouping threshold) set to 0.8 and minimal number of point samples in one density-based cluster set to 30 in MBNSFP [37]. Such parametrization allows for the successful grouping of points in whole bodies of large objects such as vehicles but simultaneously groups multiple smaller objects together. Another issue stems from points that fall below such threshold and, therefore, do not form the cluster and are treated without rigidity regularization in the methods. See an example of such DBSCAN clustering in Figure 4, where the multiple pedestrians are grouped together with a tree or a vehicle in the middle of point cloud has blue points (points not assigned to any cluster, therefore without rigidity regularization).
We conclude that such clustering parametrization leads to overfitting large objects. Therefore we designed our hard clustering approach to assign all points to clusters in a "over-segmentation" manner and gradually merge the over-segmented clusters based on flow optimization and soft-clusters.
IV-E Ablation Study
In Table V, we show metrics when adding one enhancement per time. We observe the benefits of components on the final performance on the Argoverse2 dataset. Without rigidity regularization and clusters a), the method resembles flow only to the nearest neighbors in the target point cloud and, therefore, fails to produce meaningful motions. By including the hard cluster rigidity b), we observe physically consistent motions. However, since we enforce rigidity on over-segmented clusters, the objects still deform. With the addition of soft clusters c), we are able to connect nearby points from different hard segments to rigidity terms and reject outliers, expanding the object rigidity. By guiding the hard clusterization with flows d), we acquire a small boost by connecting lonely clusters (divided by occlusions or separated by LiDAR sampling rate) back to the major rigid body.
TABLE V: Ablation study for individual method components on Argoverse2.
Module Avg. $\textit{EPE}(m)$downarrow Dyn. Fore. $\textit{EPE}(m)$downarrow Stat. Fore. $\textit{EPE}(m)$downarrow Stat. Back. $\textit{EPE}(m)$downarrow
@@ IV-E Ablation Study
IV-F Runtime Benefit of Implementing Soft Clusters as Neighborhoods
We analyze the performance-time trade-off in Figure 5 by acquiring the $EPE$ based on inference time on Argoverse1 dataset while using Neural Prior architecture with proposed Soft Clusters to demonstrate runtime difference between iterative per-cluster rigidity [37] and neighborhood parallelization ($\mathcal{L}_{soft}$). The speed was measured on NVIDIA Tesla V100 GPU on complete point clouds. We compare with optimization-based methods FastNSF [21] and MBNSF [37] with distance transform acceleration proposed by [21] and DBSCAN clustering, and Neural prior [20] without distance transform acceleration [21].
Compared to the MBNSFP [37], our method is approximately 10×10$\times$10 × faster while stopping at the same prediction error. Even though their spatial consistency regularization uses the same algorithm of computing the point-to-point displacements with outlier rejection, our method is significantly faster because we design the clusters on the same-sized neighborhoods, which can be efficiently parallelized on GPUs. In contrast to our method, MBNSFP computes eigenvectors of a matrix with variable-sized clusters [9]. Thus, MBNSFP is looping over $c$ clusters $\{N_{0},N_{1},...,N_{c}\}$ one-by-one, whereas our method allows GPU to parallel the $\mathbb{R}^{N\times k\times k}$ tensor computation for $N$ points in point cloud and $k$ neighbors simultaneously.
Figure 5: Flow estimation error vs. run time on Argoverse1 dataset. Our method in green offers a speed-accuracy trade-off that outperforms the converged baselines on the overall EPE metric.
IV-G Limitations
@@
Even though our method is reasonably fast in terms of rigidity regularization, it still does not achieve real-time performance while keeping high accuracy. Potential direction for improvement is shown in [36]. Secondly, our method does not deal with the local minima of the distance loss, resulting in sub-optimal flows even when the object is correctly clustered. These two issues are worth of the future research to improve scene flow further.
V Conclusion
@@
We presented a novel self-supervised method for 3D scene flow prediction inspired by the rigid motion of multiple objects in the scene. The method introduced overlapping cluster rigidity regularization, which achieves state-of-the-art on standard benchmarks and is tuned toward complex dynamic objects rather than static backgrounds. The framework is part of the optimization-based models family and, therefore, does not require any training data. The proposed solution adapts the rigidity mechanism to discover long tail object flows, showing that it is important, how are the points paired into the regularization mechanism.